• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 19:01:21 on June25,2004

    Entry number 126851

    keyword=Detector HV changed during run 2403

    During run 2403, we had increased the beam current to 35 uA, so we needed to change the HV on Det1 (Left arm, slot 10, ch 9 of LeCroy HV) so that we wouldn't saturate the ADC.  It was changed from -1320 V to -1270 V during this run. Gain went from ~60000 to ~40000.
    
    PS Left arm, slot 10, ch 10 was changed for a minute also (to -1220 V), but it was changed back to its nominal (-1270 V) right away.